# Computing Nonequilibrium Transport from Short-Time Transients: From Lorentz Gas to Heat Conduction in One Dimensional Chains Davide Carbone Laboratoire de Physique de l’Ecole Normale Supérieure, ENS Université PSL, CNRS, Sorbonne Université, Université de Paris, Paris, France

### v1

@@ 3.2 Anharmonic chain
As a prototypical example of an extended many-body system, we consider a one-dimensional oscillator chain with a nonlinear pinning potential [25, 24]. The system consists of $N$ particles of mass $m$, coupled by linear nearest-neighbor interactions and subject to an on-site pinning potential. The total potential energy of the system is given by
@@
It is known that the presence of pinning breaks translational invariance and suppresses momentum conservation, significantly affecting transport and relaxation properties. Heat transport is diffusive with a size-dependent thermal conductivity $\kappa(N)$ converging to a finite value in the large $N$ limit [19, 2].
@@
In the macroscopic description of continuous media, heat transport is characterized by the local heat flux density, which, according to Fourier’s law, is proportional to the temperature gradient. In this framework, the thermal conductivity $\kappa$ quantifies the material’s ability to conduct heat and is defined by the ratio between the stationary heat flux and the imposed temperature gradient. In finite one-dimensional systems subject to a temperature difference $\Delta T$ across a length $N$, this relation is commonly expressed in the form
@@ 4 Numerical Results
We next investigate the transport properties of the anharmonic chain as introduced in section 3.2. We stress that in this example we did not fully explore the computational scalability of TTCF, hence the comparison with TA in terms of computational cost will not be the focus of the present section. All simulations are performed at fixed gradient $\Delta T/N=\{0.001,1\}$ in adimensional units. In particular, the temperature of the left thermal reservoir is fixed at $T_{L}=1$, while, in the nonequilibrium regime, the temperature of the right reservoir $T_{R}$ is adjusted according to the chain length $N$ in order to maintain a constant temperature gradient. The bulk is defined as the central 70% of oscillators in the chain. The objective is to compute the thermal conductivity $\kappa$ as a function of $N$. In the linear-response regime, the pinned chain exhibits normal heat conduction, i.e., a size-independent conductivity, in agreement with previous studies [25, 2, 19].
In Fig. 4 we consider the case of smaller gradient $\Delta T/N=0.001$. In the left panel, we plot the TTCF for different values of $N$ and for an ensemble of $10^{7}$ chains. A clear spike marks the onset of the nonequilibrium phase, whose position is delayed as the chain length increases. This behavior is expected: since wave propagation in the pinned chain occurs with finite, $N$-independent speed, the time to propagate energy fluctuations between the thermostats is proportional to the chain length. Unlike simpler systems such as the Lorentz gas coupled with an isokinetic thermostat, thermal fluctuations are here much larger, even after the transient. Nevertheless, it is possible to extract the exact response after the transient, and associate to it error bars as the standard deviation of the cumulative integral considered as a stationary time series.
@@
Another feature of this nonlinear regime is that the conductivity $\kappa$ systematically decreases in the considered $N$ range, roughly as an inverse power of $N$, see right panel in Fig. 5. Since our aim here was to test the effectiveness of the method, we did not investigate this issue further. We just remark that, it is presumably due to the fact that for such large gradients heat transport may by strongly suppressed, a phenomenon described as negative differential thermal resistance [26]. Indeed, such effect is usually explained by the fact that wave spectrum depend on the energy density [38] leading to a frequency mismatch and reduced transport.
@@
To further clarify the statistical properties of our method, we compared the TTCF-based response estimate with a direct ensemble average computed from the same trajectories used for the TTCF evaluation. For the Lorentz gas, we considered up to $10^{6}$ trajectories for representative field strengths $10^{-3}$ and, a strongly nonlinear regime, $5$; for the 1D anharmonic chains, we used $10^{7}$ trajectories with corresponding temperature gradient of $10^{-3}$ and $1$. The corresponding results are reported in Figs. 7 and 8.

### v2

@@ 3.2 Anharmonic chain
As a prototypical example of an extended many-body system, we consider a one-dimensional oscillator chain with a nonlinear pinning potential [25, 24]. The system consists of NN particles of mass $m$, coupled by linear nearest-neighbor interactions and subject to an on-site pinning potential. The total potential energy of the system is given by
@@
It is known that the presence of pinning breaks translational invariance and suppresses momentum conservation, significantly affecting transport and relaxation properties. Heat transport is diffusive with a size-dependent thermal conductivity $\kappa(N)$ converging to a finite value in the large NN limit [19, 2].
@@
In the macroscopic description of continuous media, heat transport is characterized by the local heat flux density, which, according to Fourier’s law, is proportional to the temperature gradient. In this framework, the thermal conductivity $\kappa$ quantifies the material’s ability to conduct heat and is defined by the ratio between the stationary heat flux and the imposed temperature gradient. In finite one-dimensional systems subject to a temperature difference $\Delta T$ across a length NN, this relation is commonly expressed in the form
@@ 4 Numerical Results
We next investigate the transport properties of the anharmonic chain as introduced in section 3.2. We stress that in this example we did not fully explore the computational scalability of TTCF, hence the comparison with TA in terms of computational cost will not be the focus of the present section. All simulations are performed at fixed gradient $\Delta T/N=\{0.001,1\}$ in adimensional units. In particular, the temperature of the left thermal reservoir is fixed at $T_{L}=1$, while, in the nonequilibrium regime, the temperature of the right reservoir $T_{R}$ is adjusted according to the chain length NN in order to maintain a constant temperature gradient. The bulk is defined as the central 70% of oscillators in the chain. The objective is to compute the thermal conductivity $\kappa$ as a function of NN. In the linear-response regime, the pinned chain exhibits normal heat conduction, i.e., a size-independent conductivity, in agreement with previous studies [25, 2, 19].
In Fig. 4 we consider the case of smaller gradient $\Delta T/N=0.001$. In the left panel, we plot the TTCF for different values of NN and for an ensemble of $10^{7}$ chains. A clear spike marks the onset of the nonequilibrium phase, whose position is delayed as the chain length increases. This behavior is expected: since wave propagation in the pinned chain occurs with finite, NN-independent speed, the time to propagate energy fluctuations between the thermostats is proportional to the chain length. Unlike simpler systems such as the Lorentz gas coupled with an isokinetic thermostat, thermal fluctuations are here much larger, even after the transient. Nevertheless, it is possible to extract the exact response after the transient, and associate to it error bars as the standard deviation of the cumulative integral considered as a stationary time series.
@@
Another feature of this nonlinear regime is that the conductivity $\kappa$ systematically decreases in the considered NN range, roughly as an inverse power of NN, see right panel in Fig. 5. Since our aim here was to test the effectiveness of the method, we did not investigate this issue further. We just remark that, it is presumably due to the fact that for such large gradients heat transport may by strongly suppressed, a phenomenon described as negative differential thermal resistance [26]. Indeed, such effect is usually explained by the fact that wave spectrum depend on the energy density [38] leading to a frequency mismatch and reduced transport.
@@
To further clarify the statistical properties of our method, we compared the TTCF-based response estimate with a direct ensemble average computed from the same trajectories used for the TTCF evaluation. For the Lorentz gas, we considered up to $10^{6}$ trajectories for representative field strengths $10^{-3}$ and, a strongly nonlinear regime, 55; for the 1D anharmonic chains, we used $10^{7}$ trajectories with corresponding temperature gradient of $10^{-3}$ and 11. The corresponding results are reported in Figs. 7 and 8.